What is ratio of De-broglie wavelegth of proton and electron at same kinetic energy.
1
Expert's answer
2020-09-11T09:15:44-0400

solution:-

de-broglie wavelength can be written as

"\lambda=\frac{h}{p}=\frac{h}{mv}=\frac{h}{\sqrt{2mKE}}"


therefore de-broglie wavelength for proton


"\lambda _p=\frac{h}{\sqrt{2m_pKE_p}}...........eq.1"


de-brolie wavelength for electron can written as


"\lambda _e=\frac{h}{\sqrt{2m_eKE_e}}...........eq.2"



from equation 1 and 2 we can write


"\frac{\lambda _p}{\lambda _e}=\sqrt{\frac{m_e}{m_p}}"


because kinetic energy is same for proton and electron


"\frac{\lambda _p}{\lambda _e}=\sqrt{\frac{9.1\times10^{-31}}{1.67\times10^{-27}}}=\frac{0.02}{1}"


therefore ratio of de-broglie waelength of proton and electron is 0.02:1